# Soft gravitons as Goldstone modes of spontaneously broken asymptotic symmetries in de Sitter

### v1

@@ 3 Symmetry breaking: local versus global observers
$Q(\xi)=\frac{1}{2}\int d^{3}x\Pi_{\gamma}^{ij}\delta\gamma_{ij}~{}+~{}h.c.~{},$ (3.1)
where $\Pi_{\gamma}$ is the canonical momentum, was derived in [19]. Using eq.(2.3) and eq.(2.9), it was shown that the charge generating the shift in the gravitational field is equivalent to adding a soft (long) mode, such that in the shifted vacuum,
@@
[…]
@@
Following an argument similar to the standard one for $U(1)$ symmetry breaking in [31], we can now take a superposition of all the $\left|\gamma_{L}\right>$ states, which have no preference for a specific $\gamma_{L}$ value by integrating it out
$|\Psi_{\text{sym}}\rangle=\frac{1}{\mathcal{N}}\int\mathcal{D}\gamma_{ij}^{L}\psi\left[\gamma_{ij}^{L}\right]\left|\gamma_{L}\right\rangle~{}.$ (3.4)
@@
$\left\langle\Psi_{\text{sym }}\right|\gamma_{ij}(x)\left|\Psi_{\text{sym }}\right\rangle=0$ (3.5)
@@
$\psi\left[\gamma_{ij}^{L}\right]=\left\langle\gamma_{L}\mid\Psi_{\text{sym }}\right\rangle$ (3.6)
can be interpreted as the soft part of the wave function of the universe. Here, I used that the $|\gamma^{L}\rangle$ are orthogonal in the large volume limit. To understand this and what I mean by large volume limit, consider dividing a region of the space-like future infinity into $N$ patches that are uncorrelated, except for along the long-range order parameter. One could think of this as $N$ Hubble patches. The wave function of the region then factorizes into a product of the wave functions in each patch
$\Psi_{\gamma_{L}}\left(x_{1},x_{2},\ldots\right)\approx\prod_{i=1}^{N}\psi_{\gamma_{L}}\left(x_{i}\right)$ (3.7)
@@
$\left\langle\Psi_{\gamma_{L}}\mid\Psi_{\gamma_{L}^{\prime}}\right\rangle\approx\prod_{i=1}^{N}\left\langle\psi_{\gamma_{L}}\left(x_{i}\right)\mid\psi_{\gamma_{L}^{\prime}}\left(x_{i}\right)\right\rangle\propto\left(\epsilon_{\gamma_{L}{\gamma_{L}}^{\prime}}\right)^{N}\rightarrow 0\quad\text{for}\quad N\rightarrow\infty\,.$ (3.8)
@@
$\left\langle\Psi_{\gamma_{L}^{\prime}}\right|\mathcal{O}_{1}\left(x_{a}\right)\mathcal{O}_{2}\left(x_{b}\right)\cdots\mathcal{O}_{n}\left(x_{j}\right)\left|\Psi_{\gamma_{L}}\right\rangle$ $\propto$ $\left(\epsilon_{\gamma_{L}{\gamma_{L}}^{\prime}}\right)^{N-n}\prod_{i=1}^{n}\langle\mathcal{O}_{i}\rangle\rightarrow 0$ (3.9)
$\quad\text{for}\quad\gamma_{L}^{\prime}\neq\gamma_{L}\,,\quad N\rightarrow\infty\,.$
@@
The point is that if one takes the ensemble average of $\gamma_{ij}\right>$ in many local vacuum states, one will find that it vanishes. Therefore the local state is well represented by the symmetric state, $\left|\Psi_{\text{sym}}\right>$ which is a superposition of all possible values of $\gamma_{L}$. In this local state, $\gamma_{L}$ has no physical meaning, as the state does not depend on it. On the other hand, when comparing $\gamma_{ij}\right>$ in different local patches, one will see that they correlate in the infinite volume limit. The infinite volume is, therefore, well described by a single choice of $\gamma_{ij}^{L}(x)$ and the symmetry is broken by the global state, which is well represented by $\left|\Psi_{\gamma_{L}}\right>$. It is the act of measurement, or equivalently, the environment, which breaks the symmetry. Thus, once we include observers in the discussion of de Sitter, the symmetry is spontaneously broken. This emphasizes the old mantra that in any discussion of a quantum system, the observers are part of the system [32].
4 Inflation and spontaneously broken dilatation
In the case of inflation, the broken time-translation symmetry of de Sitter implies that there is a scalar curvature perturbation, $\zeta$, in addition to the usual two graviton modes. This mode corresponds to the trace part of the metric, and eq.(2.2) therefore generalizes to
$\delta\left(e^{2\zeta}\left[e^{\gamma}\right]_{ij}\right)={\cal L}_{\xi}\left(e^{2\zeta}\left[e^{\gamma}\right]_{ij}\right)~{}.$ (4.1)
From this, we deduce that at lowest order
$\delta\zeta=\frac{1}{3}\partial\cdot\xi\,.$ (4.2)
@@ 4 Inflation and spontaneously broken dilatation
$\mathcal{L}(\zeta)\propto-\dot{\zeta}^{2}+a^{-2}(\partial\zeta)^{2}$ (4.4)
which agrees with correct quadratic action of $\zeta$ after some integrations by parts [33]. This is an elegant different way of understanding why the action looks particularly simple in the comoving gauge. In inflation in 4 space-time dimensions, we therefore have 3 Goldstone modes. The two related to the spontaneously broken anisotropic scale transformation, of which the soft tensor modes, $\gamma_{ij}^{L}$ are the Goldstones, and one related to the breaking of dilations, of which the curvature perturbation, $\zeta_{L}$, is the Goldstone mode.
5 Examples
@@
5.1 Three-point consistency relation
In [19] it was shown that the explicit construction of the charge can be used to calculate the three-point function in the squeezed limit. Consider
$\left\langle 0^{\prime}\right|\gamma_{L}\gamma_{S}\gamma_{S}\left|0^{\prime}\right\rangle$ $=\langle 0|e^{-iQ}\gamma_{L}\gamma_{S}\gamma_{S}e^{iQ}|0\rangle$ (5.1)
@@ 5.1 Three-point consistency relation
$\langle 0|\left[Q,\gamma_{L}\gamma_{S}\gamma_{S}\right]|0\rangle=\left\langle\gamma_{L}\gamma_{L}\right\rangle k^{2}\frac{\partial}{\partial k^{2}}\left\langle\gamma_{S}\gamma_{S}\right\rangle$ (5.2)
@@
5.2 Infrared loop corrections
Similarly, one can also check the IR one-loop corrections of the two-point correlation function of short scalar modes during inflation can be computed by
$\left\langle 0^{\prime}\right|\zeta_{S}\zeta_{S}\left|0^{\prime}\right\rangle=\langle 0|e^{-iQ}\zeta_{S}\zeta_{S}e^{iQ}|0\rangle$ (5.3)
where an explicit evaluation and averaging over long (soft) modes reproduces the result of [18] (see [19]). Similar relations hold for tensor modes as well.
5.3 Orthogonalization of states
We showed earlier that the $\left|\gamma_{L}\right>$ states become orthogonal in the large volume limit. Since de Sitter is expanding, this is equivalent to the late time limit. We can therefore also use the results of [19] to understand the time-scale of orthogonalization of the $\left|\gamma_{L}\right>$ states. In [19] it was shown that
$\langle 0^{\prime}\mid 0\rangle=\langle 0|e^{-iQ}|0\rangle=1-\frac{1}{2}\frac{8\pi^{2}}{45}\langle\gamma_{L}\gamma_{L}\rangle+\dots$ (5.4)
Using that the variance of the soft modes in de Sitter grows as
$\langle\gamma_{L}\gamma_{L}\rangle\simeq\frac{1}{2\pi^{2}}\frac{H}{M_{p}}\log(a(t))$ (5.5)
@@ 5.3 Orthogonalization of states
$\frac{H}{M_{p}}\log(a(t))\sim\mathcal{O}(1)~{}.$ (5.6)
When using the relation $\log(a(t))=Ht$ for de Sitter, one finds that this is the equivalent of the Page time for the de Sitter, $t\sim M_{p}^{2}/H^{3}\sim R_{dS}S_{dS}$, where $R_{dS}$ is the de Sitter radius and $S_{dS}$ is the de Sitter entropy [^2] A related argument was made in the context of black holes and the information paradox already in [34]..
Thus, in de Sitter the Page time can be understood as the time scale on which we start to probe the broken phase, and no longer can characterize our state by the symmetric state.
6 Conclusions and outlook
I have shown that the soft graviton mode in de Sitter can be viewed as the Goldstone mode of the spontaneously broken asymptotic symmetry group of de Sitter. We showed that the argument holds in any dimension and generalizes straightforwardly to inflationary, quasi-de Sitter, space-times.
Through examples, I discussed how this is just another formulation of the usual consistency relations and can be viewed as the de Sitter version of the equivalence between soft-theorems and asymptotic symmetries, which, when also thought of as a gravitational memory effect [20], completes the triangle relating asymptotic symmetries, soft theorems and gravitational memory [35].
There are different applications and extensions of my work one could consider. Extending the Lagrangian of the Goldstone mode to higher orders, one has a different approach to an effective field theory of inflation or dark energy. In this context, it could also be interesting to extend our formalism to modified gravity theories or massive gravity theories. In addition, it is know that the charge, $Q$, has a dual CFT interpretation as a topological charge [36], and it would be interesting to explore the CFT dual of the spontaneous symmetry breaking. Finally, it would be interesting if a de Sitter liquid crystal, with the same symmetry-breaking pattern as discussed here, could be realized in the laboratory, as an analog system of de Sitter and eternal inflation.

### v2

@@ III Symmetry breaking: local versus global observers
$Q(\xi)=\frac{1}{2}\int d^{3}x\Pi_{\gamma}^{ij}\delta\gamma_{ij}~{}+~{}h.c.~{},$ (18)
where $\Pi_{\gamma}$ is the canonical momentum, was derived in Ferreira:2016hee . Using Eq.(3) and Eq.(9), it was shown that the charge generating the shift in the gravitational field is equivalent to adding a soft (long) mode, such that in the shifted vacuum,
@@
$0^{\prime}\right|\gamma_{ij}\left|0^{\prime}\right>$$\equiv$ $\gamma_{L}\right|\gamma_{ij}\left|\gamma_{L}\right>=\left<0\right|e^{-iQ_{{\gamma}_{L}}}\gamma_{ij}e^{iQ_{{\gamma}_{L}}}\left|0\right>$ (20)
$=$ $-i\left<0\right|\left[Q_{{\gamma}_{L}},\gamma_{ij}\right]\left|0\right>=\gamma_{ij}^{L}$
@@
Following an argument similar to the standard one for $U(1)$ symmetry breaking in Wilczek:2012jt , we can now take a superposition of all the $\left|\gamma_{L}\right>$ states, which have no preference for a specific $\gamma_{L}$ value by integrating it out
$|\Psi_{\text{sym}}\rangle=\frac{1}{\mathcal{N}}\int\mathcal{D}\gamma_{ij}^{L}\psi\left[\gamma_{ij}^{L}\right]\left|\gamma_{L}\right\rangle~{}.$ (21)
@@
$\left\langle\Psi_{\text{sym }}\right|\gamma_{ij}(x)\left|\Psi_{\text{sym }}\right\rangle=0$ (22)
@@
$\psi\left[\gamma_{ij}^{L}\right]=\left\langle\gamma_{L}\mid\Psi_{\text{sym }}\right\rangle$ (23)
can be interpreted as the soft part of the wave function of the universe. Here, I used that the $|\gamma^{L}\rangle$ are orthogonal in the large volume limit. To understand this and what I mean by large volume limit, consider dividing a region of the spacelike future infinity into $N$ patches that are uncorrelated, except along the long-range order parameter. One could think of this as $N$ Hubble patches. The wave function of the region then factorizes into a product of the wave functions in each patch
$\Psi_{\gamma_{L}}\left(x_{1},x_{2},\ldots\right)\approx\prod_{i=1}^{N}\psi_{\gamma_{L}}\left(x_{i}\right)$ (24)
@@
$\left\langle\Psi_{\gamma_{L}}\mid\Psi_{\gamma_{L}^{\prime}}\right\rangle$$\approx$ $\prod_{i=1}^{N}\left\langle\psi_{\gamma_{L}}\left(x_{i}\right)\mid\psi_{\gamma_{L}^{\prime}}\left(x_{i}\right)\right\rangle$ (25)
$\propto$ $\left(\epsilon_{\gamma_{L}{\gamma_{L}}^{\prime}}\right)^{N}\rightarrow 0\quad\text{for}\quad N\rightarrow\infty\,.$
@@
$\left\langle\Psi_{\gamma_{L}^{\prime}}\right|\mathcal{O}_{1}\left(x_{a}\right)\mathcal{O}_{2}\left(x_{b}\right)\cdots\mathcal{O}_{n}\left(x_{j}\right)\left|\Psi_{\gamma_{L}}\right\rangle$
$\propto$ $\left(\epsilon_{\gamma_{L}{\gamma_{L}}^{\prime}}\right)^{N-n}\prod_{i=1}^{n}\langle\mathcal{O}_{i}\rangle\rightarrow 0\quad\text{for}\quad\gamma_{L}^{\prime}\neq\gamma_{L}\,,\quad N\rightarrow\infty\,.$
@@
The point is that if one takes the ensemble average of $\gamma_{ij}\right>$ in many local vacuum states, one will find that it vanishes. Therefore the local state is well represented by the symmetric state, $\left|\Psi_{\text{sym}}\right>$ which is a superposition of all possible values of $\gamma_{L}$. In this local state, $\gamma_{L}$ has no physical meaning, as the state does not depend on it. On the other hand, when comparing $\gamma_{ij}\right>$ in different local patches, one will see that they correlate in the infinite volume limit. The infinite volume is, therefore, well described by a single choice of $\gamma_{ij}^{L}(x)$ and the symmetry is broken by the global state, which is well represented by $\left|\Psi_{\gamma_{L}}\right>$. It is the act of measurement, or equivalently, the environment, which breaks the symmetry. Thus, once we include observers in the discussion of de Sitter space, the symmetry is spontaneously broken. This emphasizes the old mantra that in any discussion of a quantum system, the observers are part of the system bohr .
IV Inflation and spontaneously broken dilatation
In the case of inflation, the broken time-translation symmetry of de Sitter space implies that there is a scalar curvature perturbation, $\zeta$, in addition to the usual two graviton modes. This mode corresponds to the trace part of the metric, and Eq.(2), therefore generalizes to
$\delta\left(e^{2\zeta}\left[e^{\gamma}\right]_{ij}\right)={\cal L}_{\xi}\left(e^{2\zeta}\left[e^{\gamma}\right]_{ij}\right)~{}.$ (27)
From this, we deduce that at the lowest order
$\delta\zeta=\frac{1}{3}\partial\cdot\xi\,.$ (28)
@@ IV Inflation and spontaneously broken dilatation
$\mathcal{L}(\zeta)\propto-\dot{\zeta}^{2}+a^{-2}(\partial\zeta)^{2}$ (30)
which agrees with correct quadratic action of $\zeta$ after some integrations by parts Maldacena:2002vr . This is an elegant, different way of understanding why the action looks particularly simple in the comoving gauge. In inflation in four spacetime dimensions, we therefore have three Goldstone modes. The two from the spontaneously broken anisotropic scale transformation, of which the soft tensor modes, $\gamma_{ij}^{L}$ are the Goldstones, and one from the breaking of dilations, of which the curvature perturbation, $\zeta_{L}$, is the Goldstone mode.
V Examples
@@
V.1 Three-point consistency relation
In Ferreira:2016hee it was shown that the explicit construction of the charge can be used to calculate the three-point function in the squeezed limit. Consider
$\left\langle 0^{\prime}\right|\gamma_{L}\gamma_{S}\gamma_{S}\left|0^{\prime}\right\rangle$ $=\langle 0|e^{-iQ}\gamma_{L}\gamma_{S}\gamma_{S}e^{iQ}|0\rangle$ (31)
@@ V.1 Three-point consistency relation
$\langle 0|\left[Q,\gamma_{L}\gamma_{S}\gamma_{S}\right]|0\rangle=\left\langle\gamma_{L}\gamma_{L}\right\rangle k^{2}\frac{\partial}{\partial k^{2}}\left\langle\gamma_{S}\gamma_{S}\right\rangle$ (32)
@@
V.2 Infrared loop corrections
Similarly, one can also check that the IR one-loop corrections of the two-point correlation function of short scalar modes during inflation can be computed by
$\left\langle 0^{\prime}\right|\zeta_{S}\zeta_{S}\left|0^{\prime}\right\rangle=\langle 0|e^{-iQ}\zeta_{S}\zeta_{S}e^{iQ}|0\rangle$ (33)
where an explicit evaluation and averaging over long (soft) modes reproduces the result of Giddings:2010nc (see Ferreira:2016hee ). Similar relations hold for tensor modes as well.
V.3 Orthogonalization of states
We showed earlier that the $\left|\gamma_{L}\right>$ states become orthogonal in the large volume limit. Since de Sitter spacetimes are expanding, this is equivalent to the late time limit. We can therefore also use the results of Ferreira:2016hee to understand the timescale of orthogonalization of the $\left|\gamma_{L}\right>$ states. In Ferreira:2016hee it was shown that
$\langle 0^{\prime}\mid 0\rangle=\langle 0|e^{-iQ}|0\rangle=1-\frac{1}{2}\frac{8\pi^{2}}{45}\langle\gamma_{L}\gamma_{L}\rangle+\dots$ (34)
Using that the variance of the soft modes in de Sitter spacetimes grows as
$\langle\gamma_{L}\gamma_{L}\rangle\simeq\frac{1}{2\pi^{2}}\frac{H}{M_{p}}\log(a(t))$ (35)
@@ V.3 Orthogonalization of states
$\frac{H}{M_{p}}\log(a(t))\sim\mathcal{O}(1)~{}.$ (36)
When using the relation $\log(a(t))=Ht$ for de Sitter space, one finds that this is the equivalent of the Page time for de Sitter space, $t\sim M_{p}^{2}/H^{3}\sim R_{dS}S_{dS}$, where $R_{dS}$ is the de Sitter radius and $S_{dS}$ is the de Sitter entropy Ferreira:2016hee ; Giddings:2011ze ; Giddings:2010nc . A related argument was made in the context of black holes and the information paradox already in Giddings:2007ie .
Thus, in de Sitter spacetimes, the Page time can be understood as the time scale on which we start to probe the broken phase, and can no longer characterize our state by the symmetric state.
VI Conclusions and outlook
I have shown that the soft graviton mode in de Sitter spacetimes can be viewed as the Goldstone mode of the spontaneously broken asymptotic symmetry group of de Sitter space. I showed that the argument holds in any dimension and generalizes straightforwardly to inflationary, quasi–de Sitter spacetimes.
Through examples, I discussed how this is just another formulation of the usual consistency relations and can be viewed as the de Sitter version of the equivalence between soft-theorems and asymptotic symmetries, which, when also thought of as a gravitational memory effect Ferreira:2017ogo , completes the triangle relating asymptotic symmetries, soft theorems and gravitational memory Strominger:2017zoo .
There are different applications and extensions of my work one could consider. Extending the Lagrangian of the Goldstone mode to higher orders, one has a different approach to an effective field theory of inflation or dark energy. In this context, it could also be interesting to extend our formalism to modified gravity theories or massive gravity theories. In addition, it is know that the charge, $Q$, has a dual Conformal Field Theory (CFT) interpretation as a topological charge Kehagias:2017rpe , and it would be interesting to explore the CFT dual of the spontaneous symmetry breaking. Finally, it would be interesting if a de Sitter liquid crystal, with the same symmetry-breaking pattern as discussed here, could be realized in the laboratory, as an analog system of de Sitter spacetimes and eternal inflation.